x+y+3z=-4
2x -z=-3
-x-y-2z=5

Answer by Edwin McCravy(20086)

 x + y + 3z = -4
2x     -  z = -3
-x - y - 2z =  5

Since y is already eliminated from the second equation,
eliminate y from the 1st and 2nd equations.  We only need
to add them term by term: 

 x + y + 3z = -4
-x - y - 2z =  5
----------------
          z =  1

Next we sustitute in the 2nd original equation:

     2x - z = -3
     2x - 1 = -3
         2x = -2
          x = -1

Finally substitute in the 1st original equation:

   x + y + 3z = -4
-1 + y + 3(1) = -4
  -1 + y + 3 = -4
       2 + y = -4
           y = -6

Solution is (x,y,z) = (-1,-6, 1)

Edwin